OpenStudy (anonymous):

I did this part 1)A. f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 the table n | f(n) ----------- 1 | 35 2 | ? need to find f(2) n>1 (and here n is 2, so we can use it) f(n) = f(1)+ f(n-1) replace n with 2: f(2) = f(1) + f(2-1) or f(2) = f(1)+f(1) from the table, we see f(1) is 35, so we get f(2) = 35+35 f(2) = 70 I don't know how to fill it in to find the F(3) someone please help me

OpenStudy (sidsiddhartha):

for F(3) put n=3 in the equation f(n) = f(1) + f(n-1)

OpenStudy (sidsiddhartha):

it will be \[f(3)=f(1)+f(3-1)=f(1)+f(2)\] now u have both values of f(1) and f(2) so just plug them

OpenStudy (anonymous):

can you help me more because im not good at this

OpenStudy (sidsiddhartha):

close this one and ask separately and btw welcome to open study and dont forget to appreciate the helper by giving him a medal ^_^

OpenStudy (anonymous):

I don't get it im lost how do you get that